OpenStudy (anonymous):

If a line with a slope of -2 bross the y-axis at (0,3), waht is the equation of the line. My answer was y= -2x+3 can anyone explain what i did wrong ?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Can you show me how you get that answer?

OpenStudy (anonymous):

y=mx+b slope intercept

OpenStudy (whpalmer4):

\[y = -2x\] part is correct

OpenStudy (anonymous):

Here are my choices, this is where i got a little confused y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3

OpenStudy (anonymous):

the y intercept is where the line crosses through the right axis which was pinpointed at postivite 0,3 which is 3 as a whole number

OpenStudy (31356):

Here is an example to refer to: http://www.algebra.com/cgi-bin/Equations.faq.question.84744.html

OpenStudy (whpalmer4):

\(y = -2x+3\) does go through the point \((0,3)\). \[3 = -2(0)+3\] Ah, maybe you screwed up when converting to one of those answer… \[y = -2x+3\]If we add \(2x\) to both sides:\[y+2x = -2x + 2x + 3\]\[y+2x = 3\]

OpenStudy (anonymous):

y axis*

OpenStudy (whpalmer4):

y-intercept is just the value of y when x = 0. it's just a number when referring to it in slope-intercept form (b), otherwise it is usually written as \((0,b)\)

OpenStudy (anonymous):

ohhh ok sorry if i'm missing a couple of steps my online class requires me to teach myself lol i guess i forgot a couple of things while scanning down my notes ,thanks whpalmer4 huge help

OpenStudy (whpalmer4):

you're welcome. give me a shout if you ever need help understanding something, and I'll try to help you out.

OpenStudy (anonymous):

Sure thing bro thanks!